Solve the system: x + y - z = 2
2x + y + z = 3
3x + y + z = 5

Answer by Alan3354(69443)

x + y - z = 2
2x + y + z = 3
3x + y + z = 5
--------------- Subtract #3 from #2
-x = -2
x = 2
-------
x + y - z = 2 --> y - z = 0
2x + y + z = 3 --> y + z = -1
----------------------
Add those 2
2y = -1
y = -1/2
------------
Sub into eqn #1:
2 -1/2 - z = 2
z = -1/2